Two particles of the same mass are projected from the same place with the same velocity u, such that their ranged are same. If h1 and h2 are the maximum heights attained by them then the relation between h1,h2 and R will be:
a) R= h1 / h22
b) R2=  h1  / h2
c)  R2=  h1 h2
d)  R2= 16 h1 h2

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Range is same for angles of projection is 90
R = (u^2 sin^2 theta)/(g) h1 = (u^2 cos^2 theta)/(2 g)
and h2 = (u^2 sin^2 (90 - theta))/(2 g)
= (u^2 cos^2 theta)/(2 g)
Hence 
(sqrt(h1 h2)) = (u^2 sin theta cos theta)/(2 g) = (1)/(4) [(u^2 sin 2theta)/(g)] = (R)/(4)
So , R = 4sqrt(h1h2) option  D is correct 
Good Luck